### v1

@@ 6 Related Work
Abductive Reasoning with LLMs.
Our work relates to LLM-based investigative agents that perform multi-step reasoning over structured environments Yao et al. (2023); Bei et al. (2025). These systems follow a reason-retrieve-synthesize loop and have been applied to software diagnosis Roy et al. (2024); Xu et al. (2025), cloud operations Chen et al. (2024a); Tian et al. (2025), and graph-augmented QA Sun et al. (2024); Luo et al. (2024). Recent work has also begun to introduce belief-centric abstractions to stabilize LLM agent reasoning by making intermediate assumptions explicit in language Lidayan et al. (2025); Kim et al. (2024); He et al. (2025). These approaches improve consistency by guiding how beliefs are formed or queried, but they treat beliefs as prompt-level artifacts rather than first-class, mutable state. While effective for episodic investigations, they assume monotonic and global updates, and lack explicit mechanisms for belief revision when new evidence contradicts earlier hypotheses.
@@
System Graph as Mutable State Hypotheses Explicit Update Semantics Revision Trigger Propagation Controller Invariants Data Dynamicity Graph Scale
ReAct Yao et al. (2023) append global episodic small
Graph-RAG Edge et al. (2024); Sun et al. (2024) $\circ$ append global summ. episodic medium
Tool-assisted RCA Roy et al. (2024); Xu et al. (2025); Zhang et al. (2025b) $\circ$ $\circ$ append global $\circ$ streaming large
Multi-agent (STRATUS) Chen et al. (2025) $\circ$ $\circ$ append global $\circ$ streaming large
GALA Tian et al. (2025) $\circ$ $\circ$ append global/hybrid $\circ$ streaming large
Ours (EoG) revise msg passing streaming unbounded
Legend: ✓ = yes, ✗ = no, $\circ$ = partial/implicit.
Graph-RAG and Graph-Augmented Agents.
Graph-RAG systems Edge et al. (2024); He et al. (2024); Liu et al. (2025b) extend retrieval to knowledge graphs, enabling multi-hop reasoning and improved coverage over large corpora. Recent work explores optimizing the interface between knowledge graphs and LLMs Marković et al. (2025); Coppolillo (2025) and integrating knowledge graph reasoning into agent workflows Besta et al. (2025). GraphRAG Edge et al. (2024) constructs entity knowledge graphs, partitions them into community hierarchies, and pre-generates summaries aggregated at query time. Think-on-Graph Sun et al. (2024) performs iterative reasoning by exploring knowledge graph relations, while Reasoning on Graphs Luo et al. (2024) generates faithful reasoning plans grounded in graph structure. Theorem-of-Thought Abdaljalil et al. (2025) runs three specialized agents in parallel, stores each agent’s step-wise reasoning as a graph, and then applies Bayesian belief propagation to score these graphs and select the most coherent one. StructGPT Jiang et al. (2023) provides a general framework for reasoning over structured data. However, their inference semantics are append-only: evidence is accumulated globally, with no explicit hypothesis representation, revision trigger, or localized propagation operator. This makes them order-sensitive and brittle under evolving evidence, especially when graphs are large and cannot be fully materialized in context.
Diagnostic Agents in Operations.
LLM agents for diagnosis in cloud operations span several architectures: tool-assisted RCA agents Roy et al. (2024); Xu et al. (2025); Zhang et al. (2025b); Deshpande et al. (2025), multi-agent systems like STRATUS Chen et al. (2025), and graph-augmented workflows like GALA Tian et al. (2025). Recent work also provides systematic taxonomies of failure modes in such diagnostic agents Ma et al. (2025). These systems are explicitly abductive: they generate explanatory hypotheses from partial evidence. Graph-based systems like Groot Wang et al. (2021) and causal approaches like RADICE Tonon et al. (2025) use service dependency graphs to localize faults; similar approaches have been applied to microservice systems Zhu et al. (2024) and IoT environments with dynamic causal graphs Tang et al. (2025). GraphLocator Liu et al. (2025a)-style traversal is captured by our Deterministic Controller’s graph-guided expansion, and GALA/SynergyRCA-style Xiang et al. (2025) evidence aggregation is subsumed by the Context Packet; neither exposes explicit belief revision or localized message passing. Nevertheless, all these systems rely on implicit hypotheses, global summarization, and workflow control rather than controller-enforced invariants over mutable graph state. Even graph-aware workflows remain monotonic, without first-class revision semantics or message passing that localizes updates to affected neighborhoods. Our ReAct baseline subsumes the inference semantics of these approaches: it has access to all diagnostic tools (topology queries, log/metric analysis, trace inspection) plus shell utilities and code execution, strictly more capability than published RCA agents. Yet even the best-performing ReAct configuration (Gemini 3 Flash) achieves only 74% majority@3 F1 on ITBench, compared to 92% for EoG with the same model. This 18-point gap persists across model families, confirming that gains arise from state semantics (explicit beliefs, revision, propagation), not domain engineering or tool access.
Causal Inference.
Causal discovery and effect estimation Wan et al. (2024); Vowels et al. (2022); Yao et al. (2021) target different outputs (DAGs, treatment effects) under static, fully-observed data; they do not model revisable explanatory hypotheses under streaming partial evidence and are thus not comparable to abductive graph reasoning for baseline consideration.
Putting EoG in perspective.
@@
Our contribution is an inference abstraction for abductive reasoning under streaming evidence and unbounded graphs, where global context is infeasible and explanations must be revised online. The baselines above cover the strongest LLM instantiations of investigative reasoning; differences arise from state semantics, not domain or prompting. Table 5 summarizes the comparison across inference-semantic axes. We compare against representative ReAct, Graph-RAG, and diagnostic agents, and ablate revision and propagation—demonstrating that gains arise from state semantics rather than domain-specific engineering.
@@ 7 Conclusion
We framed operational diagnosis as abductive inference over graphs and introduced Semantic Belief Propagation, which separates local LLM reasoning from deterministic control to enable non-monotonic belief revision. Experiments on ITBench show gains in accuracy, stability, and token efficiency over ReAct-style agents. This architecture yields reproducible, auditable explanations under realistic production constraints.
@@ B.5 Initialization Prompts
The investigation begins with a multi-phase initialization that acquires bounded context and selects initial candidates.
@@ B.8 Implementation Notes
Appendix C LLM-as-a-Judge Evaluation Framework
@@
C.1 Entity Mapping Challenge
@@
C.1.1 Ground Truth Structure
@@ C.1.4 The Mapping Challenge
C.2 Judge System Prompt
@@
LAAJ System Prompt You are an expert AI evaluator specializing in Root Cause Analysis (RCA) for complex software systems. Your role is to act as a fair, consistent, and precise judge when evaluating RCA model outputs. You will be provided with: 1. A Ground Truth (GT) JSON object: true root cause(s), failure propagation path(s), and entity definitions. 2. A Generated Response JSON object: output from an RCA model to evaluate. Phase 1: Normalization and Entity Mapping
Before any scoring, normalize and map entities from Generated Response to Ground Truth based on explicit evidence from entity metadata. Do not infer mappings based on position in propagation chain.
Definition of a Confident Match:
A Generated Response entity is a confident match ONLY IF its name field clearly corresponds to the filter and kind of the GT entity.
Entity Name Format:
Generated Response entities use format: namespace/Kind/name
Examples: otel-demo/Deployment/flagd, otel-demo/Service/frontend
CRITICAL - Alias Handling:
If a predicted entity matches an ALIAS of a root cause entity, it MUST be counted as correct. The GT.aliases field contains arrays of equivalent entity IDs. If ANY entity in an alias group has root_cause: true, then matching ANY entity in that alias group counts as correctly identifying the root cause.
@@ C.3 Evaluation Metrics
C.3.1 Root Cause Entity Identification
@@
C.3.2 Root Cause Reasoning Accuracy
@@
C.4 Evaluation Prompt Template

### v2

@@ 6 Related Work
ReAct Yao et al. (2023)
Graph-RAG Edge et al. (2024)
RCA Roy et al. (2024)
STRATUS Chen et al. (2025)
GALA Tian et al. (2025)
EoG (Ours)
Graph as mutable state $\circ$ $\circ$ $\circ$ $\circ$
Hypotheses explicit $\circ$ $\circ$ $\circ$
Belief revision
Propagation glob g.sum glob glob hyb msg
Controller invariants $\circ$ $\circ$ $\circ$
Data dynamicity epis epis str str str str
✓=yes, ✗=no, $\circ$=partial. epis=episodic, str=streaming.
Propagation: how updates spread—glob=entire context passed each step, g.sum=global with summarization, hyb=mix of global/local, msg=local message passing.
Our work relates to LLM-based investigative agents performing multi-step reasoning over structured environments Yao et al. (2023); Bei et al. (2025); Roy et al. (2024); Xu et al. (2025); Chen et al. (2024a); Tian et al. (2025); Sun et al. (2024); Luo et al. (2024). Recent work introduces belief-centric abstractions Lidayan et al. (2025); Kim et al. (2024); He et al. (2025), but treats beliefs as prompt-level artifacts rather than mutable state. These systems assume monotonic updates and lack explicit mechanisms for belief revision in presence of contradictory evidence. Causal discovery and effect estimation Wan et al. (2024); Vowels et al. (2022); Yao et al. (2021) while related, target different outputs under static, fully-observed data and do not model revisable explanatory hypotheses under streaming regimes.
Graph-RAG systems Edge et al. (2024); He et al. (2024); Liu and others (2025); Marković et al. (2025); Coppolillo (2025); Besta et al. (2025); Sun et al. (2024); Luo et al. (2024); Jiang et al. (2023) enable multi-hop reasoning over knowledge graphs. Theorem-of-Thought Abdaljalil et al. (2025) uses Bayesian belief propagation over parallel agent reasoning graphs to aggregate multiple reasoning attempts; in contrast, EoG uses message passing to correct sequential mistakes as new evidence emerges.
Diagnostic agents for cloud operations Roy et al. (2024); Xu et al. (2025); Zhang et al. (2025b); Deshpande et al. (2025); Chen et al. (2025); Tian et al. (2025); Ma et al. (2025) generate explanatory hypotheses from partial evidence. Graph-based systems Wang et al. (2021); Tonon et al. (2025); Zhu et al. (2024); Tang et al. (2025); Liu et al. (2025); Xiang et al. (2025) use dependency graphs to localize faults. All these rely on append-only inference without explicit revision or localized propagation, making them brittle under evolving evidence in unbounded graphs.
@@
Table 5 summarizes the comparison of EOG with representative ReAct, Graph-RAG, and diagnostic agents, and ablate revision and propagation, demonstrating that gains arise from state semantics rather than domain-specific engineering.
@@ 7 Conclusion
Diagnostic investigations over massive, heterogeneous data with hidden dependency structure require consistent correctness; stochastic success is insufficient when verification carries risk. We introduced EoG, a disaggregated architecture that separates local abductive reasoning via LLMs from deterministic graph traversal and belief propagation, achieving $7\times$ higher Majority@k F1 on ITBench scenarios. This separation of concerns generalizes to other domains where explanations must be derived from environments with inherent graphical structure.
8 Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.4 SRE Domain Tools (MCP Server)
Operational Graph Construction.
The build_topology tool constructs the operational graph $G$ by combining static information (application architecture, service dependencies declared in configuration) with dynamic runtime state (e.g., pod-to-node bindings, replica counts, endpoint mappings from Kubernetes). Such operational graphs are readily available through commercial observability platforms—including IBM Instana IBM (2025), Dynatrace Dynatrace (2025), and Datadog Datadog (2025)—which automatically discover and maintain service topologies through agent instrumentation and API introspection.
@@ B.5 Initialization Prompts
The investigation begins with a multi-phase initialization that sets the algorithm variables from §4.3: select_investigation_window determines time window $W$; filter_events and filter_spec_changes acquire bounded context; bootstrap selects initial candidates $V_{\mathcal{O}}$ for the ActiveSet.
@@ B.8 Implementation Notes
B.9 Properties and Correctness
B.9.1 Minimality and Correctness
The algorithm seeks a Frontier $\mathcal{F}$ of irreducible origins.
$\mathcal{F}=\{v\in V_{S}:$ $L_{v}=\textsc{Origin}\land\nexists u\in V_{S}$ (1)
$\text{s.t. }L_{u}=\textsc{Origin}\land u\rightsquigarrow_{E_{S}}v\}$
By definition, $\mathcal{F}$ is minimal relative to the discovered graph $G_{S}$. If no node is labeled Origin (i.e., $\mathcal{F}=\emptyset$), the system falls back to best-effort ranking of candidates based on accumulated evidence. The correctness relies on three assumptions:
  1. 1.
    Completeness: Relevant data exists and is retrievable via CxC.
  2. 2.
  3. 3.
    Connectivity: Origins are reachable from anomalous entities $V_{\mathcal{O}}$ via the topology or discovered edges.
If these hold, the message passing ensures that evidence from origins drives the observed symptoms, correcting any initial misclassifications caused by the exploration order.
On the role of the deterministic controller.
The Soundness assumption merits clarification: EoG does not eliminate LLM stochasticity—belief assignments remain LLM-dependent. The deterministic controller ensures systematic exploration, belief bookkeeping, and propagation logic, but does not guarantee that local inferences are correct. The architecture’s value lies in bounding the scope of each LLM decision to local context and enabling revision when contradictory evidence emerges via message passing. Reliability gains (Majority@$k$ improvements) arise from this structured exploration and revision mechanism, not from eliminating stochasticity in individual inferences.
B.9.2 Termination
B.9.3 Handling Inconclusive Results
The agent may not always produce a definitive diagnosis. Two failure modes are possible: (1) evidence absence—the relevant observability data is missing; and (2) evidence blindness—the policy $\pi_{\text{abd}}$ lacks the domain knowledge to recognize available evidence as diagnostically relevant. In either case, nodes are labeled Defer. If no Origin is identified at termination (i.e., frontier $\mathcal{F}=\emptyset$), the Controller falls back to a best-effort ranking of candidate nodes based on accumulated evidence, explicitly signaling uncertainty to the operator.
Appendix C Analyzing Controller Failures
The ReAct paradigm conflates reasoning (what to investigate) with execution (how to invoke tools). This overloading of the LLM leads to measurable controller instability. We analyze three failure modes—plan abandonment, tool repetition, and syntactic failures—that motivate the separation of concerns in EoG.
(a) Plan abandonment rate.
(b) Tool repetition rate.
(c) Tool failure categories.
Plan abandonment.
Tool repetition.
Syntactic failures.
Figure 7(c) categorizes tool invocation failures. A substantial fraction are strictly syntactic—malformed arguments, type errors, invalid parameter combinations—rather than semantic reasoning failures Chen et al. (2024b). These errors indicate that LLMs are suboptimal drivers for precise tool orchestration when simultaneously reasoning about complex domains. In contrast, conventional non-deterministic programs tend to excel in tool invocation and execution when they have access to the right contextual information.
C.1 MAST Failure Mode Taxonomy
We use MAST (Multi-Agent System Trajectories) UC Berkeley (2025) to systematically categorize controller failures in our agent trajectories. MAST provides a structured taxonomy that classifies failures across three inter-agent conversation stages: Pre-Execution, Execution, and Post-Execution.
Figure 8: MAST failure mode taxonomy. Failures are categorized into three main categories: System Design Issues (44.2%)—failures stemming from how the agent system is architected; Inter-Agent Misalignment (32.3%)—failures in communication and coordination; and Task Verification (23.5%)—failures in validating task completion.
Figure 8 shows the taxonomy with prevalence rates across our trajectories:
System Design Issues (44.2%).
The largest category encompasses architectural failures:
  • 1.1 Disobey Task Specification (11.8%): Agent ignores explicit task instructions.
  • 1.2 Disobey Role Specification (1.5%): Agent acts outside its designated role.
  • 1.3 Step Repetition (15.7%): Agent repeats previously executed steps without progress.
  • 1.4 Loss of Conversation History (2.8%): Agent loses track of prior context.
  • 1.5 Unaware of Termination Conditions (12.4%): Agent fails to recognize when task is complete.
Inter-Agent Misalignment (32.3%).
Communication and coordination failures:
  • 2.1 Conversation Reset (2.2%): Agent restarts conversation inappropriately.
  • 2.2 Fail to Ask for Clarification (6.8%): Agent proceeds despite ambiguity.
  • 2.3 Task Derailment (7.4%): Agent deviates from the assigned task.
  • 2.4 Information Withholding (0.8%): Agent fails to share relevant information.
  • 2.5 Ignored Other Agent’s Input (1.9%): Agent disregards feedback.
  • 2.6 Reasoning-Action Mismatch (13.2%): Agent’s stated reasoning contradicts its actions.
Task Verification (23.5%).
Failures in validating task completion:
  • 3.1 Premature Termination (6.2%): Agent stops before task completion.
  • 3.2 No or Incomplete Verification (8.2%): Agent fails to verify its output.
  • 3.3 Incorrect Verification (9.1%): Agent incorrectly validates erroneous output.
Figure 9: Failure mode counts across models. GPT-OSS-120B and Kimi-K2 exhibit higher failure counts across most categories compared to Gemini-3-Flash, with Step Repetition (1.3) and Reasoning-Action Mismatch (2.6) being the most prevalent failure modes.
Figure 9 compares failure mode prevalence across three models. Key observations: (1) Step Repetition (1.3) is the most common failure across all models, with GPT-OSS-120B showing 86 instances versus 36 for Gemini-3-Flash. (2) Reasoning-Action Mismatch (2.6) shows similar patterns, indicating a fundamental limitation of the ReAct paradigm where LLMs must simultaneously plan and execute. (3) Gemini-3-Flash exhibits consistently lower failure counts, correlating with its higher diagnostic accuracy.
Figure 10: Failure mode distribution in successful vs. failed traces for Gemini-3-Flash. Green bars show prevalence in successful traces; red bars show prevalence in failed traces.
Figure 11: Failure mode distribution in successful vs. failed traces for Kimi-K2. Green bars show prevalence in successful traces; red bars show prevalence in failed traces.
Figure 12: Failure mode distribution in successful vs. failed traces for GPT-OSS-120B. Green bars show prevalence in successful traces; red bars show prevalence in failed traces.
Figures 1012 reveal which failure modes differentiate successful from failed traces:
  • Discriminative failures: Incorrect Verification (3.3) shows the largest gap between failed and successful traces across all models (66% vs 14% for Gemini-3-Flash, 83% vs 43% for Kimi-K2, 94% vs 50% for GPT-OSS-120B). This indicates that verification failures are strongly predictive of task failure.
  • Non-discriminative failures: Step Repetition (1.3) appears at high rates in both successful and failed traces (71% vs 59% for Gemini-3-Flash), suggesting that some repetition may be a natural part of iterative exploration rather than a direct cause of failure.
  • Model-specific patterns: GPT-OSS-120B shows elevated failure rates across nearly all categories in failed traces, while Gemini-3-Flash maintains lower baseline rates even in failed cases.
These findings motivate the core architectural decision in EoG: by separating deterministic control (state management, termination detection, verification) from non-deterministic reasoning (causal inference), we eliminate the System Design Issues (44.2%) and Task Verification failures (23.5%) that account for over two-thirds of observed failures.
Appendix D LLM-as-a-Judge Evaluation Framework
@@
Evaluation cost considerations.
Each agent run produces a structured diagnosis that must be evaluated against ground truth with semantic matching (entity aliasing, reasoning quality assessment). The LAAJ evaluation requires an LLM call per (scenario, run) pair, incurring non-trivial cost at scale. For our benchmark of 50 scenarios with $k=5$ runs per model, evaluating a single model requires 250 LAAJ calls. Due to this cost, we focus our detailed analysis on a representative subset of frontier models (Gemini 2.5 Pro/Flash, Claude Opus 4.5, o4-mini) that span the capability spectrum, while reporting aggregate metrics for additional models.
D.1 Entity Mapping Challenge
@@
D.1.1 Ground Truth Structure
@@ D.1.4 The Mapping Challenge
D.2 Judge System Prompt
@@ D.3 Evaluation Metrics
D.3.1 Root Cause Entity Identification
@@
D.3.2 Root Cause Reasoning Accuracy
@@
D.4 Evaluation Prompt Template